The mass of the Earth is 6.45c10^24kg and it's radius is 3.38x1010. compute the density of the earth using power of ten notation
1
Expert's answer
2019-08-05T12:01:40-0400
"\rho=\frac{m}{V}=\frac{m}{\frac{4\pi r^3}{3}}"

"\rho =\frac{3m}{4\pi r^3}"

"\rho =\frac{3(6.45\cdot 10^{24})}{4\pi (3.38\cdot 10^{10})^3}=3.99\cdot 10^{-8}\frac{kg}{m^3}"